OpenStudy (anonymous):

the slope of a line is -1/2 the point (6,-7) is on the line. determine the equation of this line

OpenStudy (solomonzelman):

you know that slope intercept form is y=mx+b where m is the slope an d b is the y intercept. slope= -1/2 so ----> y=-1/2x+b to find be plug in your point (6,-7).

OpenStudy (anonymous):

what is the answer though